· Let e be the event that a computer crashed and O the event that a computer overheated Suppose we have the following probabilities: P(C) 0.4 P(Cno) 0.25 P(O) 0.3 P(C no) 0.05 )-0.15 P(C no) 0.55 Answer the questions below. Round your answers below to 2 decimals. (a) Find the probability that the computer crashed, given that the computer overheated. (b) Find the probability that the computer crashed, given that the computer did not (c) Find the probability that the computer did not overheat, given that the computer did not crash. (d) Are the events of a computer crashing and a computer overheating independent events? Explain 2. A random sample of 300 adults are classified below by gender and their level of education attained Education Male Female Elementary 5 Secondary 6 College 37 12 13 If a person is picked at randon from this group, find the probability that a) the person is a female. given that the person has a college degre (b) the person does not have only au elementary exdueation. given that the person is a

1&2
Answer #1

Solution:-

2)

a) The person is a female, given that the person has a college degree is 0.5059.

Female P(FemaCollege) College P( P College

P(College) = 85/300

P(College) = 0.28333

P(Female and college) = 43/300

P(Female and college) = 0.14333

pFemale0.14333 College0.28333 P(

Female0.5059 College

b) Notelementary0.6519 P( Male

Motelementary. Male P(Not.elementarnmale) P(male) P(

P(Male) = 158/300

P(Male) = 0.5267

P(Not elementary and Male) = 103/300

P(Not elementary and Male) = 0.34333

P(Notelementary 0.34333 Male 0.5267

Notelementary0.6519 P( Male